OpenStudy (anonymous):

find the dirivitive by limit process 2x^2+x+1

OpenStudy (anonymous):

you can't find a limit of x^2 ?

OpenStudy (anonymous):

\[f(x)=2x^2+x+1~~\Rightarrow~~f'(x)=\lim_{h\to0}\frac{\left(2(x+h)^2+(x+h)+1\right)-\left(2x^2+x+1\right)}{h}\] \[f'(x)=\lim_{h\to0}\frac{2x^2+4xh+2h^2+x+h+1-2x^2-x-1}{h}\] \[f'(x)=\lim_{h\to0}\frac{4xh+2h^2+h}{h}=\cdots\]